OpenStudy (anonymous):

will give medal and become a fan to best answer. solve 0.6(y + 3) = 4.8 A. 5 B. 3 C. 13 D. 1.8

OpenStudy (anonymous):

a

OpenStudy (anonymous):

0.6y+1.8=4.8 4.8-1.8=3 3/0.6=?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0.6*5=3

OpenStudy (anonymous):

so it is B?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

it was right! thanks!

OpenStudy (anonymous):

because... when you simplify your problem you have 0.6y+1.8=4.8 you subtract 1.8 from 4.8 and you get 3 3 divided by 0.6 is 5

OpenStudy (anonymous):

you are welcome